40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1678     Entry time: Tue Jun 16 14:02:16 2009
Author: rob 
Type: Update 
Category: Locking 
Subject: locked 

The IFO is locked, at the operating point (zero CARM offset).  The problem with reducing the residual CARM offset in the last stage appears to have been because the common mode gain was getting too high, and so the loop was going unstable at high frequencies. 

The cm_step script is currently a confusing mess, with all the debugging statements.  I'll clean it up this afternoon and check that it still works.

ELOG V3.1.3-